OpenStudy (angelic):

A water ballon has a balloon has a diameter of 6 inches. how many cubic inches of water does it contain when it is 3/4 full

OpenStudy (anonymous):

V = 4/3πr^3 V = 4/3π216 V = 2(216)π/3 V = 144π Now, it's 3/4 full. 144π(3)/4 V = 108π or V ≈ 339.29

OpenStudy (anonymous):

Find the volume of a sphere whose diameter is 6, the sphere contains that much water when it is full. So when it is 3/4 full it contains 3/4 * the volume you just calculated